OpenStudy (anonymous):

WIll give medal and fan!!! What is the maximum or minimum value of the function? What is the range? y = –2x2 + 32x –12 (1 point) maximum: 116 range: y 116 maximum: –116 range: y –116 maximum: 116 range: y 116 maximum: –116 range: y –116

OpenStudy (solomonzelman):

this is an opening down parabola, and we know that it has no absolute minimum, since it will be going infinitely down on both ends.

OpenStudy (anonymous):

Yes

OpenStudy (solomonzelman):

you can find the vertex of this, and this will be the absolute maximum of the function. need help findning the vertex?

OpenStudy (anonymous):

Yes.

OpenStudy (solomonzelman):

\(\large\color{black}{ y = -2x^2 + 32x -12 }\) first go ahead an factor out of \(\large\color{black}{ -2 }\) for me please.

OpenStudy (anonymous):

Okay!

OpenStudy (solomonzelman):

what do you get, after doing this?

OpenStudy (solomonzelman):

are you offering me a calculator to fill in values so that I can factor it? I know how to factor it:) I want to see though how you would factor it out of -2.

OpenStudy (anonymous):

take the derivative: -4x+32 and set it equal to zero. -4x+32=0 -4x=-32 x=8 Take second derivative: -4 which indicates the curve is concave down. Indicated it is a max. Maximum 116 at x=8 Range: (-inf, 116)

OpenStudy (anonymous):

but I don't know what the open would be?

OpenStudy (solomonzelman):

you don't need the derivative for an absolute maximum of a parabola.

OpenStudy (anonymous):

i dont get it.

OpenStudy (solomonzelman):

you know since the coefficient of x^2 is negative that it is going down. This means that it will have the absolute maximum, but not absolute minimum. you can just find the vertex of it

OpenStudy (solomonzelman):

the vertex will be the absolute maximum.

OpenStudy (solomonzelman):

\(\large\color{black}{ y = -2x^2 + 32x -12 }\) please factor out of -2 for me, (you should know how to do it, if you are already mentioning derivatives)

OpenStudy (solomonzelman):

you derivatives are correct, but there is no need in doing this.

OpenStudy (anonymous):

a = -2x^2

OpenStudy (solomonzelman):

a =-2x^2 ? what does this mean?

OpenStudy (anonymous):

I am sorry I don't know how to do it :(((

OpenStudy (solomonzelman):

if you can't factor the \(\large\color{black}{ y = -2x^2 + 32x -12 }\) out of -2 (on the right hand side of the equation) then I don't know how to do. You want to tell me you are capable of finding the derivative and of applying it to find the absolute maximum, but can't simply factor?

OpenStudy (anonymous):

I forgot hwo tooo. ]